Today Kerry read 7 less than twice the number of pages she read yesterday. If she read a total of 122 pages, how many did she read yesterday?
Answer by Cromlix(4381)

Hi there,
Make the number of pages read yesterday = 'x'
Today Kerry would have read 2x - 7
Over the two days she will have read:-
x + 2x - 7 = 122
3x - 7 = 122
3x = 122 + 7
3x = 129
x = 129/3
x = 43 pages
Kerry read 43 pages yesterday.
Hope this helps :-)